# Position Coupling: Leveraging Task Structure for Improved Length Generalization of Transformers

### v1

@@ Abstract
Even for simple arithmetic tasks like integer addition, it is challenging for Transformers to generalize to longer sequences than those encountered during training. To tackle this problem, we propose position coupling, a simple yet effective method that directly embeds the structure of the tasks into the positional encoding of a (decoder-only) Transformer. Taking a departure from the vanilla absolute position mechanism assigning unique position IDs to each of the tokens, we assign the same position IDs to two or more “relevant” tokens; for integer addition tasks, we regard digits of the same significance as in the same position. On the empirical side, we show that with the proposed position coupling, a small (1-layer) Transformer trained on 1 to 30-digit additions can generalize up to 200-digit additions (6.67$\times$ of the trained length). On the theoretical side, we prove that a 1-layer Transformer with coupled positions can solve the addition task involving exponentially many digits, whereas any 1-layer Transformer without positional information cannot entirely solve it. We also demonstrate that position coupling can be applied to other algorithmic tasks such as addition with multiple summands, $N\times 2$ multiplication, copy/reverse, and a two-dimensional task. [^1] Our codebase is available at github.com/HanseulJo/position-coupling.
@@ 1 Introduction
Since the appearance of a sequence-to-sequence deep neural architecture called Transformer (Vaswani et al., 2017), it has brought tremendous success in various fields including natural language process (NLP) (Thoppilan et al., 2022; Chowdhery et al., 2023; Gemini et al., 2023; OpenAI, 2023) and many applications such as mathematical reasoning and theorem proving (Lewkowycz et al., 2022; Wu et al., 2022; Trinh et al., 2024). Despite its triumph, it has recently been illuminated that Transformers often lack the ability of length generalization (Anil et al., 2022; Deletang et al., 2023; Zhang et al., 2023; Press et al., 2022). It refers to a special kind of out-of-distribution generalization capability to extrapolate the model’s performance to longer sequences than those encountered during training. Understanding length generalization is of great importance because the lack of length generalization provides evidence that language models do not genuinely understand the structure of a given task. Improving Transformer’s length generalization has received much attention, particularly because the time/memory complexities for training Transformers grow up to quadratically in the sequence length.
@@
Figure 1: Methods for length generalization in the integer addition task. We report exact-match accuracies (markers: medians over experiments; light area: 95% confidence intervals). We employ the reversed format and zero-paddings (Lee et al., 2024) into the input sequence. With position coupling, we achieve more than 95% exact-match (EM) accuracy for up to 200-digit additions with Transformers trained on up to 30-digit additions. For index hinting (Zhou et al., 2024a), we separately test absolute positional embedding (APE) with a random starting position ID (mimicking the original implementation by Zhou et al. (2024a)) and without positional encoding (NoPE) (Kazemnejad et al., 2023) (as tested by Zhou et al. (2024b)).
In this paper, we propose position coupling, a simple yet effective method for length generalization that directly embeds the structure of the tasks into a Transformer. In contrast to the vanilla absolute position mechanism assigning unique and consecutive position IDs to each token, we assign the same position IDs to certain input tokens that are semantically relevant. Coupling such tokens together helps the model learn to solve the task regardless of the length of the given input sequence. For example, in the addition task, it is important to consider the significance of digits, so we couple the positions at the same significance.
@@ 1.1 Summary of Contributions
  • With position coupling, we achieve a robust and near-perfect generalization up to 200-digit additions by training Transformers on up to 30-digit additions, which is a 6.67×6.67$\times$6.67 × extrapolation of the operand lengths (Figure 1, Section 4). It is promising since it was not clear whether the length generalization on the addition task can be solved reliably with Transformers (Zhou et al., 2024b).
@@
  • We empirically show that position coupling can effectively address various tasks beyond addition, including addition with multiple summands (Section 6.1), multiplication between $N$-digit and 2-digit integers (Section 6.2), and copy/reverse operations (Section 7.1). In particular, we provide a theoretical construction of a 2-layer Transformer that solves the multiplication between $N$-digit and 2-digit integers for exponentially large $N$ (Theorem 6.1). We also verify that position coupling can aid Transformers in learning tasks with multi-dimensional structures (Section 7.2).
@@ 2 Preliminaries
We focus on decoder-only Transformers that solve the tasks using next-token prediction. Since we study deterministic tasks with a unique answer, we consider greedy decoding throughout the paper.
2.1 Next-token Prediction with Decoder-only Transformers
Figure 2: Schematic of solving an integer addition task instance using next-token prediction with a decoder-only Transformers. Here we assume a basic input format (plain, no zero-padding), which is different from that we used in our experiments.
@@ 2.2 Data Formats
Each task in this work is represented as a collection of sequences of the form ‘(query)=(response)’: given a query, our task is to infer the response correctly. Thus, we only care about the result of the next-token prediction for the ‘=’ token and the tokens in the response (except its last token). That is, we only compute the losses and accuracies for those output tokens, as depicted in Figure 2.
@@
Wrapping with BOS/EOS token(s).   It is conventional in NLP to put BOS/EOS (beginning-/end-of-sequence) tokens at the beginning/end of the sequence. Lee et al. (2024) use the same token ‘\$’ for BOS and EOS tokens and observe that it is beneficial to wrap each sequence with the \$ token when solving the addition task. We do not observe any significant difference in the performance between sequences with the same and different BOS/EOS tokens.
@@ 3 Position Coupling: A Method for Length Generalization
We propose position coupling, which assigns position IDs that directly encode the structure of given tasks. Here, we explain the general position ID assignment rule of position coupling in two steps and then move on to the specific example of addition.
@@
Next, for each group of tokens, we assign a sequence of consecutive numbers (usually, positive integers) as position IDs, starting from a random number (at training time) or a fixed predetermined number (at evaluation time). We use random position IDs at training time for inducing length generalization by enabling all position embedding vectors to be trained, up to a pre-defined hyperparameter of maximum position ID (max_pos).222 It explicitly determines the maximum testable length of sequence that a transformer can handle. In the case of the decimal integer addition task, the maximum possible generalizable length is ${\tt max\_pos}-2$. Models trained with larger ${\tt max\_pos}$ can handle longer sequences; however, training becomes more difficult as the frequency of each position ID appearing during training decreases. Very importantly, we assign the same position IDs to the tokens in all groups that are relevant to each other for solving the given task: we refer to this as “coupling the positions”. Lastly, we set 0 as the position IDs of special tokens like BOS/EOS tokens and the PAD token (padding for minibatch training and evaluation).
@@ 3.1 Position Coupling for Decimal Integer Addition Task
Figure 3: Position coupling for decimal integer addition task, showing $653+49=702$ with appropriate input formats. The starting position ID ‘6’ is an arbitrarily chosen number.
@@
We illustrate position coupling for the decimal integer addition task (or addition task for short). To study the length generalization of the addition task, we regard each digit (0–9) as a single token. For illustration, if we need a concrete example, we will use the addition $653+49=702$.
@@
Before applying the position coupling, we adopt an input format similar to Lee et al. (2024) so that we reverse the response, but we use zero-padding and wrapping with BOS/EOS token ‘\$currency-dollar\\$\$’ at the same time. For example, ‘$653+49=702$’ becomes ’\$$\$653+049=2070\$$’.
@@
We partition the tokens in the sequence into three groups: (1) first operand & ‘$+$’, (2) second operand, and (3) ‘$=$’ & response (the sum). Then each number token is “unique” in the corresponding group in terms of significance, which naturally induces a one-to-one correspondence between (most of) the tokens across different groups. We group ‘$=$’ and the sum together because these tokens are where we perform next-token prediction.
@@
Now we assign the coupled position IDs to the tokens. Let us say that the random starting number is 6. We assign 6, 7, and 8 to the tokens in the operands, and assign 5, 6, 7, and 8 to the tokens in the sum in a reversed order: see Figure 3. We remark that, first, we assign 9 as position IDs of ‘$+$’ and ‘$=$’ tokens because they are adjacent to the number token with position ID 8, even if there are no ‘significances’ for those tokens. Second, we assign 5 as a position ID of the most significant digit of the sum (which may be ‘0’ due to the zero-padding) just because it is next to the number token with position ID 6, even though there are no other corresponding tokens in the other groups (operands). We also note that we do not group the ‘$+$’ token with the second operand to prevent unnecessary coupling between ‘$+$’ and the most significant digit of the sum (position ID 5).
@@
Remark.   A concurrent work by McLeish et al. (2024) proposes an analogous approach in the context of solving arithmetic tasks, while they employ a different input format; they flip both the query and the response but do not apply any paddings. A more detailed comparison with our work is provided in Section 8.
3.2 Comparison with Index Hinting
@@
Even though the idea of implanting the structure of a task into the positional encoding is novel, there is an existing approach named index hinting (Zhou et al., 2024a) that applies a similar idea but to the input sequence. Index hinting is an input augmentation technique that places position markers in front of the tokens to couple the semantically relevant tokens. For example, Zhou et al. (2024a) transform ‘$653+49=702$’ into ‘${\tt a}0{\tt b}6{\tt c}5{\tt d}3+{\tt a}0{\tt b}0{\tt c}4{\tt d}9={\tt a}0{\tt b}7{\tt c}0{\tt d}2$’ with some zero-paddings, where ${\tt a}$, ${\tt b}$, ${\tt c}$, and ${\tt d}$ are consecutive index hints. Here, the starting hint character ${\tt a}$ is randomly selected during training, similar to our method of choosing the starting position ID. The reversed format and BOS/EOS tokens can be applied as well.
@@
One way in which index hinting differs from position coupling is that it doubles the input sequence length. This is because the position information and the token information do not merge: the index hints and the normal tokens are mapped to separate token embedding vectors which are alternately placed in the input embedding matrix. As a result, a Transformer must figure out the correspondence between each adjacent pair of an index hint and a normal token. Moreover, the doubled input length can require up to four times the training time and memory. In contrast, position coupling explicitly combines token and position information: every token embedding and corresponding position embedding are mixed into a single vector. Hence, a Transformer can effortlessly utilize the positional structure of the task, without hurting the training time. We highlight that, as will be mentioned in Section 4.1, position coupling exhibits better length generalization than index hinting.
@@
4 Experiments on Addition Task
@@
Model and Training.   As baseline models, we train 1-layer, 4-head, decoder-only Transformers (with roughly 4M trainable parameters) from scratch. We set max_pos as 202 so that the maximum testable length of summands is 200. We do not use packing or shifting for simplicity of implementation. Since we manually put coupled position IDs with a random starting index during training, we can train all the position embeddings without packing and shifting.
@@ 4.1 Results
Longer Training Sequences Lead to Longer Generalizable Lengths.   We train separate models with $D_{\max}\in\{10,20,30,40\}$ and evaluate the models on (up to) 200-digit additions. For each run of training, we choose and evaluate the best model in terms of the validation loss for 200-digit additions. We run each experiment 8 times with 2 different random seeds for data generation and 4 different random seeds for model initialization & stochastic optimization, unless mentioned otherwise. We summarize all hyperparameters in Table 1 of Appendix B. The result is showcased in Figure 4. We decide that a model successfully generalizes to a certain length of operands (referred to as “generalizable length”) if the median EM accuracy exceeds 95%.
We observe that the generalizable length becomes longer as we train on longer training sequences. The generalizable length is 70 for the models trained on additions involving 1–10 digits, 135 for models trained on 1–20, and 200 for 1–30 and 1–40. We believe that we could achieve even longer generalizable length for the models trained on 1–40 if a larger max_pos is used with careful training techniques.
@@
Figure 4: Ablation on the trained operand lengths.
A prior work by Zhou et al. (2024b) provides a similar analysis on the addition tasks. Combining appropriate input format and advanced PE, they achieve $\geq$98EM accuracy for 100-digit additions with a 6-layer 8-head model trained on 1–40. Moreover, they achieve a generalizable length of 45 for a model trained on 1–30, 25 for 1–20, and 10 for 1–10 (no length generalization). One big difference between their analysis and ours is they report the maximum accuracy for each testing length over trials, while we report the medians. Thus, we choose a bit lower threshold (95%) for generalizability than theirs. To better the comparison, we also report the maximum accuracies in Section A.1.
Shallower Models Generalize Better.   Since Zhou et al. (2024a) and Zhou et al. (2024b) choose 6-layer 8-head models as their base models, we also test our method to deeper models. The experimental details can be found in Table 2 of Appendix B. The results (with models trained on 1–40 digits) displayed in Figure 5 show that, even if deeper models can generalize up to some extent, the performance deteriorates as the model gets deeper. We attribute this phenomenon to the problem of optimization. Since the theoretical construction of a 1-layer addition Transformer (that will appear in Section 5.1) naturally extends to larger architectures, deeper models have at least as much generalization capability as shallower models. Thus, the difficulty of training deep neural networks must be the reason for the degrading performance.
The performance of the 1-layer 8-head model is a bit worse than that of our base model (shown in Figure 4), which can be attributed to the reduction of expressive power due to inversely-proportional shrink of the dimension per attention head (Bhojanapalli et al., 2020).
@@
Figure 5: Ablation on the number of layers.
@@
Comparison with Index Hinting.   We test index hinting by running the code we implemented ourselves since the original code is unavailable. From Figure 1, we observe that index hinting indeed helps the model to length-generalize more than the baselines (NoPE & random-start APE). However, the generalizable lengths of the models trained with index hinting do not extend further than 50; the models completely fail starting from the length 70. We also observe that enough amount of depth is required for Transformers with index hinting to achieve high enough training and in-distribution validation accuracies. Particularly, the training accuracies of 1-layer models do not deviate from near zero. Thus, we only present the results for the 6-layer 8-head model as done by Zhou et al. (2024a).
Scaling the Generalizable Lengths Up to 500.   We demonstrate the scalability of our proposed position coupling approach for large lengths of up to 500. Specifically, we again train a depth-1 decoder-only model for the addition task and evaluate the performance for instances with up to 500 digits. The results are shown in Figure 6. We notice that at a train length of 160, we achieve excellent length generalization for 500-digit addition. On the other hand, training on sequences of length up to 40 or 80 is insufficient for extreme length generalization. While these results lead to a smaller length generalization ratio of $500/160=3.125$, as opposed to the ratio of $6.67$ achieved when training on sequences of length up to $30$, the results demonstrate that position coupling, as an approach, is highly scalable. The experimental details can be found in Table 3 of Appendix B.
Figure 6: The full sequence accuracy obtained by training a depth-1 decoder-only model on the addition task. We see that while training with sequences of length up to 40 and 80 is insufficient for generalization to large lengths, at training length 160 we achieve strong performance for lengths up to 500.

### v2

# Position Coupling: Improving Length Generalization of Arithmetic Transformers Using Task Structure

@@ Abstract
Even for simple arithmetic tasks like integer addition, it is challenging for Transformers to generalize to longer sequences than those encountered during training. To tackle this problem, we propose position coupling, a simple yet effective method that directly embeds the structure of the tasks into the positional encoding of a (decoder-only) Transformer. Taking a departure from the vanilla absolute position mechanism assigning unique position IDs to each of the tokens, we assign the same position IDs to two or more “relevant” tokens; for integer addition tasks, we regard digits of the same significance as in the same position. On the empirical side, we show that with the proposed position coupling, our models trained on 1 to 30-digit additions can generalize up to 200-digit additions (6.67$\times$ of the trained length). On the theoretical side, we prove that a 1-layer Transformer with coupled positions can solve the addition task involving exponentially many digits, whereas any 1-layer Transformer without positional information cannot entirely solve it. We also demonstrate that position coupling can be applied to other algorithmic tasks such as $N\times 2$ multiplication and a two-dimensional task. Our codebase is available at github.com/HanseulJo/position-coupling.
Figure 1: Methods for Length Generalization in the Integer Addition Task.   We report exact-match (EM) accuracies (markers: medians over experiments; light area: 95% confidence intervals). We employ the reversed format and zero-paddings (Lee et al., 2024) into the input sequence. With our proposed position coupling, we achieve more than 95% exact-match accuracy for up to 200-digit additions with decoder-only Transformers trained on up to 30-digit additions. For index hinting (Zhou et al., 2024a), we separately test absolute positional embedding (APE) with a random starting position ID (mimicking the original implementation by Zhou et al. (2024a)) and without positional encoding (NoPE) (Kazemnejad et al., 2023) (as tested by Zhou et al. (2024b)).
@@ 1 Introduction
Since the appearance of a sequence-to-sequence deep neural architecture called Transformer (Vaswani et al., 2017), it has brought tremendous success in various fields including natural language process (NLP) (Thoppilan et al., 2022; Chowdhery et al., 2023; Gemini et al., 2023; OpenAI, 2023) and many applications such as mathematical reasoning and theorem proving (Lewkowycz et al., 2022; Wu et al., 2022; Trinh et al., 2024). Despite its triumph, it has recently been illuminated that Transformers often lack the ability of length generalization (Anil et al., 2022; Deletang et al., 2023; Zhang et al., 2023; Press et al., 2022). It refers to a special kind of out-of-distribution generalization capability to extrapolate the model’s performance to longer sequences than those encountered during training. Understanding length generalization is of great importance because a lack of it provides evidence that language models do not genuinely understand the structure of a given task. Improving Transformer’s length generalization has received much attention, particularly because the time/memory complexities for training Transformers grow up to quadratically in the sequence length.
@@
In this paper, we propose position coupling, a simple yet effective method for length generalization that directly embeds the structure of the tasks into a Transformer. In contrast to the vanilla absolute position mechanism assigning unique and consecutive position IDs to each token, we assign the same position IDs to certain input tokens that are semantically relevant. Coupling such tokens together helps the model learn to solve the task regardless of the length of the given input sequence. For example, in the addition task, it is important to consider the significance of digits, so we couple the positions at the same significance (unique in each operand and the answer).
@@ 1.1 Summary of Contributions
  • With position coupling, we achieve a robust and near-perfect generalization up to 200-digit additions by training Transformers on up to 30-digit additions, which is a 6.67×6.67$\times$6.67 × extrapolation of the operand lengths (Figure 1, Section 4). It is promising since it was unclear whether the length generalization on the addition task can be solved reliably with Transformers (Zhou et al., 2024b).
@@
  • We empirically demonstrate that position coupling can effectively address various tasks beyond addition, including multiplication between $N$-digit and 2-digit integers (Section 6.1, in which we also provide a theoretical construction of a 2-layer Transformer that solves this task for exponentially large $N$). We also verify that position coupling can aid Transformers in learning tasks with multi-dimensional structures (Section 6.2). Moreover, we evaluate position coupling on some other tasks (addition with multiple operands, copy/reverse) in Appendix B.
@@ 2 Preliminaries
We focus on decoder-only Transformers that solve the tasks using next-token prediction (See Appendix A for a brief background on it). Since we study deterministic tasks with a unique answer, we consider greedy decoding throughout the paper.
@@ 2.1 Data Formats
Each task in this work is represented as a collection of sequences of the form ‘(query)=(response)’: given a query, our task is to infer the response correctly. Thus, we only care about the result of the next-token prediction for the ‘=’ token and the tokens in the response (except its last token). That is, we only compute the losses and accuracies for those output tokens.
@@
Wrapping with BOS/EOS token(s).   It is conventional in NLP to put BOS/EOS (beginning-/end-of-sequence) tokens at the beginning/end of the sequence. Lee et al. (2024) use the same token ‘\$’ for BOS and EOS tokens and observe that it is beneficial to wrap each sequence with the \$ token when solving the addition task. We do not observe any significant difference in the performance between sequences with the same and different BOS and EOS tokens.
@@ 3 Position Coupling: A Method for Length Generalization
We propose position coupling, which assigns position IDs that directly encode the structure of given tasks. Here, we explain the general position ID assignment rule of position coupling in two steps.
@@
Next, for each group of tokens, we assign a sequence of consecutive numbers (usually, positive integers) as position IDs, starting from a random number (at training time) or a fixed predetermined number (at evaluation time). We use random position IDs at training time for inducing length generalization by enabling all position embedding vectors to be trained, up to a pre-defined hyperparameter of maximum position ID (max_pos).111 The hyperparameter max_pos determines the maximum testable sequence length that a Transformer can handle. Note that the idea of random assignment of position IDs is similar to randomized PE (Ruoss et al., 2023) although it is different since it assigns a sequence of increasing integers which are generally not consecutive. Very importantly, we assign the same position IDs to the tokens in all groups that are relevant to each other for solving the given task: we refer to this as “coupling the positions”. Lastly, we set 0 as the position IDs of special tokens like BOS/EOS tokens and the PAD token (padding for minibatch training and evaluation).
@@ 3.1 Position Coupling for Decimal Integer Addition Task
Figure 2: Position coupling for decimal integer addition task, displaying $653+49=702$ with appropriate input formats. The starting position ID ‘6’ is an arbitrarily chosen number.
@@
We illustrate position coupling for the decimal integer addition task (or addition task for short). To study the length generalization of the addition task, we regard each digit (0–9) as a single token. We will use an explicit example of the addition ‘$653+49=702$’ for illustration.
@@
Before applying the position coupling, we adopt an input format similar to Lee et al. (2024) so that we reverse the response, but we use zero-padding and wrapping with BOS/EOS token ‘\$currency-dollar\\$\$’ at the same time. For example, ‘$653+49=702$’ becomes ‘\$$\$653+049=2070\$$’.
@@
We partition the tokens in the sequence into three groups: (1) first operand & ‘$+$’, (2) second operand, and (3) ‘$=$’ & response (which we call ‘sum’). Then each number token is “unique” in the corresponding group in terms of significance, which naturally induces a one-to-one correspondence between (most of) the tokens across different groups. We group ‘$=$’ and the sum together because these tokens are where we perform next-token prediction.
@@
Now we assign the coupled position IDs to the tokens. Most importantly, we assign the same position ID to the digits of the same significance. Let us say that the random starting number is 6. In our example, we assign 6, 7, and 8 to the tokens in the operands, and assign 5, 6, 7, and 8 to the tokens in the sum in a reversed order: see Figure 2. We remark that, first, we assign 9 as position IDs of ‘$+$’ and ‘$=$’ tokens because they are adjacent to the number token with position ID 8, even if there are no ‘significances’ for those tokens. Second, we assign 5 as a position ID of the most significant digit of the sum (which may be ‘0’ due to the zero-padding) just because it is next to the number token with position ID 6, even though there are no other corresponding tokens in the other groups (operands). We also note that the ‘$+$’ token is not grouped with the second operand and is not given the ID 5; this is to prevent unnecessary coupling between ‘$+$’ and the most significant digit of the sum.
@@
Remark.   A concurrent work by McLeish et al. (2024) proposes an analogous approach for solving arithmetic tasks, while they employ a different input format. We provide a detailed comparison with our work in Appendix A.
@@
Comparison with Index Hinting.   Even though the idea of implanting the structure of a task into the positional encoding is novel, there is an existing approach named index hinting (Zhou et al., 2024a) that applies a similar idea but to the input sequence. Index hinting is an input augmentation technique that places position markers in front of the tokens to couple the semantically relevant tokens. For example, Zhou et al. (2024a) transform ‘$653+49=702$’ into ‘${\tt a}0{\tt b}6{\tt c}5{\tt d}3+{\tt a}0{\tt b}0{\tt c}4{\tt d}9={\tt a}0{\tt b}7{\tt c}0{\tt d}2$’ with some zero-paddings, where ${\tt a}$, ${\tt b}$, ${\tt c}$, and ${\tt d}$ are consecutive index hints. Here, the starting hint character ${\tt a}$ is randomly selected during training, similar to our method of choosing the starting position ID. The reversed format and BOS/EOS tokens can be applied as well.
@@
One way in which index hinting differs from position coupling is that it doubles the input sequence length. This is because the position information and the token information do not merge: the index hints and the normal tokens are mapped to separate token embedding vectors which are alternately placed in the input embedding matrix. As a result, a Transformer must figure out the correspondence between each adjacent pair of an index hint and a normal token. Moreover, the doubled input length requires up to 4×4$\times$4 × the training time and memory consumption. In contrast, position coupling explicitly combines token and position information: every token embedding and corresponding position embedding are mixed into a single vector. Hence, a Transformer can effortlessly utilize the positional structure of the task, without hurting the training time. We highlight that, as will be mentioned in Section 4.1, position coupling exhibits better length generalization than index hinting.
@@
4 Experiments on the Addition Task
@@
Model and Training.   We train decoder-only Transformer models from scratch. We properly choose max_pos so that the maximum testable length of summands is 200. We do not use packing or shifting for simplicity of implementation. Since we manually put coupled position IDs with a random starting index during training, we can train all the positions without packing and shifting. We run each experiment 8 times with 2 different random seeds for data generation and 4 different random seeds for model initialization & stochastic optimization unless mentioned otherwise. We summarize all hyperparameters in Appendix C.
@@ 4.1 Results
Figure 3: Ablation on the trained operand lengths (1-layer 4-head models).
Longer Trained Sequences Lead to Longer Generalizable Lengths.   We train 1-layer 4-head models with $D_{\max}\in\{10,20,30,40\}$ and evaluate them on up to 200-digit additions.For each run of training, we choose and evaluate the best model in terms of the validation loss for 200-digit additions. The result is showcased in Figure 3. We decide that a model successfully generalizes to a certain length of operands (referred to as “generalizable length”) if the median EM accuracy exceeds 95%.
We observe that the generalizable length becomes longer as we train on longer training sequences. The generalizable length is 70 for the models trained on additions involving 1–10 digits, 135 for models trained on 1–20, and 200 for 1–30 and 1–40. We believe that we could achieve even longer generalizable length for the models trained on 1–40 if we use a larger max_pos. We note that we could scale up the generalizable length to 500 by training with lengths 1–160: refer to Section B.1. Although each test sample contains the operands of the same length, we also provide an extended evaluation on test samples with operands of different lengths: see Section B.2.
@@
Figure 4: Ablation on the number of layers (trained with position coupling).
Ablation on Number of Layers.   Since Zhou et al. (2024a) and Zhou et al. (2024b) choose 6-layer 8-head models as their base models, we also test our method to deeper models. The results evaluated with models trained on 1–30 digits are displayed in Figure 4, whose experimental details are listed in Table 2. Overall, the performances are well extrapolated to test lengths (longer than trained lengths) and are similar for the models with 1–5 layers. For the 6-layer model, however, the performance slightly deteriorates. We hypothesize that the performance degradation is due to the bad implicit bias of deeper models (learning shortcuts only to achieve in-distribution generalization) when learning a simple algorithm to solve the task. Since the theoretical construction of a 1-layer addition Transformer (that will appear in Section 5.1) naturally extends to larger architectures, deeper models have at least as much generalization capability as shallower models. We believe exploring a theoretical explanation for the bad implicit bias of large models on low-complexity tasks is a promising research direction. We also highlight that we present median accuracies over multiple runs, while Zhou et al. (2024b) report maximum accuracies. To better the comparison, we also report the maximum accuracies (for the experiments in Figures 3 and 4) in Section B.3, showing that our 6-layer models can achieve near-perfect generalization for up to 200-digit additions as well.
Figure 5: Ablation on the data formats (1-layer 4-head models trained with position coupling).
Figure 6: Ablation on the data formats (6-layer 8-head models trained with position coupling).
Ablation on Data Formats.   Our input format is primarily selected to simplify the algorithm of solving the addition task, not through extensive ablation studies. Thus, we are not arguing that our choice of input format is empirically optimal for training Transformers. However, since the data format is one of the crucial factors in general machine learning, we test various formats for 1-layer 4-head (Figure 5) and 6-layer 8-head models (Figure 6). The results clearly show that the performance varies with input formats, as they affect the complexity of the algorithm that the model should learn.
Small models (1-layer 4-head) achieve near-perfect generalization when the numbers are zero-padded and the response or all the numbers are reversed. We believe this is because the combination of zero-padding and reversing enabled a small Transformer to learn a simple length-generalizing algorithm. Zero-padding seems crucial since it aids length generalization to some extent even without reversing. Without reversing any numbers, however, even the in-distribution performance slightly decays.
Larger models (6-layer 8-head) perform better than small models when the numbers are no longer zero-padded or reversed. We believe this is because the task-solving algorithm without reversing or zero-padding that the model should learn is more sophisticated, which larger models can learn more easily. Contrarily, we observe a slight degradation in performance when we add zero-padding and reversing in the larger model, which suggests that the model may have learned a “shortcut” due to its (overly) strong expressive power relative to the problem’s complexity.
@@
Comparison with Index Hinting.   We test index hinting by running the code we implemented ourselves since the original code is unavailable. From Figure 1, we observe that index hinting indeed helps the model to length-generalize more than the baselines (NoPE & random-start APE). However, the generalizable lengths of the models trained with index hinting do not extend further than 50; the models completely fail starting from the length 70. We also observe that Transformers with index hinting require enough depth to achieve high enough training and in-distribution validation accuracies. Particularly, the training accuracies of 1-layer models do not deviate from near zero. Thus, we only present the results for the 6-layer 8-head model as done by Zhou et al. (2024a).
Comparison & Combination with RoPE.   We also examine the possibility of combining position coupling and RoPE (Su et al., 2024): See Section B.4 for the experimental results and details.